OpenStudy (anonymous):

If Dawn's lawn has two areas that are 10ft x 8ft,another area 22 ft x 4 ft,another 6 ft x 15 ft and the last area is 23 ft x 5 ft, what is the square footage of her lawn? If the seed is spread at a rate of 5lbs per 1,000ft^2,and each bag is 2lbs,how many bags will she need to buy?

OpenStudy (amistre64):

2(10*8) + 22(4) + 15(6) + 23(5) = total area of the yard. total area/5lbs = x lbs per 1 area. the last parts I see, but cant right put my finger on it yet :)

OpenStudy (amistre64):

(total area/1000) times 5 should be the amount of lbs of seed she needs

OpenStudy (amistre64):

once you know the total amount of seed required, figure out how many 2lb bags will do it ;)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

So,.45385= 2,265......then what?

OpenStudy (amistre64):

....45385 = which part?

OpenStudy (anonymous):

.453 multiply by 5 gave me 2,265 I'm still not sure how many bags of seeds

OpenStudy (amistre64):

2(10*8) + 22(4) + 15(6) + 23(5) 160 + 88 + 80 + 115 = 160443 right?

OpenStudy (amistre64):

443...calculator lied to me lol

OpenStudy (anonymous):

I have 453

OpenStudy (amistre64):

your right, 15(6) = 90...tried that in my head

OpenStudy (amistre64):

453/1000 = .0453 right? i tried that in my head too :)

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

so she has less than 1000 ft^2 acres, so she needs less then 5lb of seeds...

OpenStudy (amistre64):

she has 4.53% of 1000 ft^2 so I would assume she only needs 4.53% of 5 lbs of seeds.... let me check that...

OpenStudy (anonymous):

It's asking how many bags will she need to buy.

OpenStudy (amistre64):

you got any multiple choice answers to pick from?

OpenStudy (anonymous):

nope

OpenStudy (amistre64):

you got that 1000 ft^2 right? or is that spose to be 100ft^2 ?

OpenStudy (anonymous):

The first part is right 1000

OpenStudy (amistre64):

the total square footage of her yard has to be the sum of the parts..... and we did that right :) 10(8) + 10(8) + 22(4) + 6(15) + 23(5)....cant change that

OpenStudy (anonymous):

Well,you said that it's 5 lbs that she'll need ...so each bag being 2 lbs,then would that be 21/2 bags?

OpenStudy (amistre64):

it says she uses 5 lb seeds to do 1000 ft^2....she only has 453 ft^2 right?

OpenStudy (amistre64):

so for every ft^2 she uses how many lbs of seed? 5 x ------ = ----- should be the ratio 1000 1

OpenStudy (amistre64):

or: 5 x ------ = ----- should be an equal ratio as well 1000 453 5(453)/1000 = 2.265 lbs

OpenStudy (anonymous):

Yes,I have those numbers but still don't know the number of bags of seeds to buy.

OpenStudy (amistre64):

she needs 1 bag and a a little more..so if she buys 2 bags total, she should be able to do it.

OpenStudy (amistre64):

or 1.1325 bags if theyll let her ;)

OpenStudy (anonymous):

So do you think that the answer should be 2 or 1 and a half?

OpenStudy (anonymous):

hahaha! That's funny!

OpenStudy (amistre64):

if they want an integer answer, I would say 2; if they want a precise anwer, I would go with 1.1325 bags

OpenStudy (anonymous):

O.k.,that's helpful.Thanks!